OpenStudy (anonymous):

Bruce is 36 years old. son is 8 years old . in how many years will bruce be three times as old a his son?

OpenStudy (anonymous):

@LynFran , @Crissy15 and , @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

36+6=42 8+6=14 42/14=3 6 years

OpenStudy (anonymous):

6

OpenStudy (anonymous):

where'd you get the original 6 from?

OpenStudy (anonymous):

thx